# Dual-Domain Homogeneous Fusion with Cross-Modal Mamba and Progressive Decoder for 3D Object Detection

### v1

@@ III-B Homogeneous Voxel Fusion Network
Figure 4: Pipeline of Mamba-based Voxel Fusion. We utilize the bidirectional Mamba block, as illustrated in the right figure, to process both single-modal and multi-modal voxel features. Within the 3D U-Net architecture, the merge-and-split operation seamlessly switches between intra-modal and cross-modal Mamba blocks, facilitating global alignment and fusion of multi-modal information.
Figure 5: Details of Intra-Modal and Cross-Modal BEV Mamba. We apply a 4-directional cross-scan to unfold the BEV feature into 1D sequences, which helps construct a more comprehensive spatial relationship. In the cross-modal Mamba block, the parameters for the SSMs of each modality are jointly computed from the Image and LiDAR BEV features, while $Z_{I}$ and $Z_{L}$ are used to adaptively modulate the weights of their respective modalities.
@@
Where $\bigtriangleup$ denotes the time-scale parameter and $A$ denotes the learnable matrix. $B$, $C$, and $\bigtriangleup$ are inferred from the input sequence $x_{i}$ through linear layers, ensuring the input-specific of the SSM model. To ensure complete perception, both forward and backward SSMs are implemented. Their output sequences are added together and modulated by $Y$ to produce the final results.
@@
In CV-Mamba, we directly combine multi-modal voxels in the unified 3D space. It’s important to note that the height range of image voxels $Z_{I}$ is twice that of LiDAR voxels $Z_{L}$. Therefore, we need to multiple the z indices of LiDAR voxels by two before combination. Similar to IV-mamba, these voxels are unfolded into 1D sequences and fed into the SSM. While some voxels may occupy the same position, we treat them as distinct elements in the sequence rather than merging them as in [25]. According to the theoretical proof in [34], the forward-backward Mamba is a special version of a global transformer, enabling unaligned features to match their corresponding ones in the other modality. Then, the output voxels with different modalities are separated and sent back to their perspective branches, where they are downsampled by sparse 3D convolution. Notably, in most existing works, fusing features with different distributions[23, 25, 22, 42] requires operations such as nearest-neighbor search or grouping, while our designed elegant global merge-and-split operation makes the fusion process more efficient and straightforward. Alternating between IV-Mamba and CV-Mamba effectively bridges the modality gap and alleviates the negative impact of deteriorated data from each modality on the fused features.
@@
Finally, Inverse sparse convolutions are used to upsample voxels features back to the original scale. The output features $V_{I}$ and $V_{L}$ are not directly combined to avoid interfere brought by distinctive voxel distributions. Instead, they are separately fed into the homogeneous BEV fusion network and the multimodal feature mixing module in the decoder.
III-C Homogeneous BEV Fusion Network
@@
In Homogeneous BEV Fusion Network, the features from the homogeneous voxel network, $V_{I}^{’}$ and $V_{L}^{’}$, are transformed into BEV grids through sparse height compression [43] — specifically, max-pooling is applied in voxel features belonging to the same BEV pixel, yielding $B_{I}^{{}^{\prime}}$ and $B_{L}^{{}^{\prime}}$. They are then concatenated with $B_{I}$ and $B_{L}$ to inject more 3D semantic and structure details into BEV space. It is important to note that during SAFS, some potential foreground voxels may be filtered out due to low semantic or depth scores. The LSS compensates for this by retaining image information of the entire environment. In addition, both SAFS and LSS share the same depth map. Since LSS has not undergone feature filtering, it preserves all the backpropagated gradients, which helps accelerate the convergence of depth estimation during training.
@@
As shown in Fig. 4, we also applied the Mamba-based fusion paradigm, which includes both intra-modal and cross-modal BEV Mamba (IB-Mamba and CB-Mamba) modules to generate high-quality BEV features. First the concatenated features into the IB-Mamba, enabling the network to globally perceive modality-specific information. Following the SS2D block proposed in [34], we apply four-directional cross scanning to unfold the image, which helps construct a comprehensive spatial relationship. Then, as described in Equations 2 and 3, the SSM is performed simultaneously across all four directions. After passing through LayerNorm, the outputs are summed and modulated by $Y$ for final results.
@@
In CB-Mamba, to further align multi-modal features in the dense BEV space and adaptively fuse them, we extend the SS2D operation originally designed for single-modal tasks to cross-modal fusion. As shown in Fig. 4, the parameters for the SSM computation are generated from the concatenated tensor, guiding the global rectification for both image and LiDAR features. We describe the SSM parameterization using the following formulas.
@@
Where, the superscripts $1-4$ denote the parameters employed for the four-directional scanning in SS2D. Furthermore, the modulation parameters $Y_{I}$ and $Y_{L}$ are also generated by concatenating features and are used to reweight multi-modal features:
@@
It enables the network to autonomously favor the more reliable modality, thereby improves adaptability in complex scenarios.
@@
After CB-Mamba, the output BEV tensor is fed into the BEV backbone for further feature propagation, yielding $B_{out}$, which is then utilized as the input for query generation.
@@
Figure 6: Details of Hard Instance Activation. First, self-attention among easy queries is used to extract spatial relationships between objects. Then, query-to-BEV attention is applied to activate latent hard instances within the BEV features.
III-D Progressive Query Generation
@@
In DDHFusion, the BEV branch is used for efficient generation of target candidates. While many previous works have attempted to achieve this in the range view[44, 45] or sparse 3D space[46, 47], the former is impacted by object occlusions, and the latter often relies on center-voting strategies, which are sensitive to uneven voxel or point densities. In contrast generating queries in the BEV domain is relatively simple. First, we generate a heatmap from $B_{out}^{’}$ with convolution. As outlined in [14], queries can be obtained by processing response scores using non-maximum suppression based on 3x3 max pooling, followed by the top-k selection. However, some hard targets, characterized by corrupted information such as sparse surface points or blurred image textures. After height compression, these features may be further deteriorated. Moreover, since targets are typically small in the BEV plane, their features may be overshadowed during convolution operations in the BEV backbone.
@@
To address this, [48] propose a multi-stage query generation strategy. It groups queries into different difficulty levels and employs residual blocks to activate BEV features between levels to focus on harder instances. In this work, we simplify the approach by dividing queries into two groups: easy queries $Q_{e}$ and hard queries $Q_{h}$. The corresponding two-stage extraction process is detailed in following equations:
${H_{e}}=Heatmaphead({B_{out}})$ (8)
${Q_{e}}=Topk(NMS({B_{out}}))$ (9)
$M=I-Maxpooling(H)$ (10)
$H^{{}^{\prime}}=Heatmaphead(Resnet({B_{out}}))$ (11)
$Q_{h}=Topk(NMS(M\bullet{H^{{}^{\prime}}}))$ (12)
@@
Where $M$ represents the mask generated from $Q_{e}$, which is used to obtain $Q_{h}$ by masking the regions corresponding to $Q_{e}$. When training, we compute the heatmap loss of the second stage only within the regions of $M$, guiding the network to focus on hard targets. However, activating BEV feature by the residual block has limited effectiveness—it can only aggregate local information from difficult regions to enhance responses. In the driving environment, the spatial relationships between objects exhibit certain regularities, such as the placement of vehicles and barriers. Based on this observation, we propose a hard instance activation (HIA) module, which leverages easy queries $Q_{e}$ to activate features in difficult regions. Specifically, we modify Equation 11 as follows:
${H^{{}^{\prime}}}=Heatmaphead(HIA({\color[rgb]{1,0,0}Q_{e}},{B_{out}}))$ (13)
@@
As shown in Fig 6, in HIA, we first embed categories – in form of one-hot vectors – and positions into the easy queries. Then, we utilize self-attention to extract relationships between objects, resulting in enhanced queries $Q_{e}^{{}^{\prime}}$. Then, $B_{out}$ is flattened and regarded as query and $Q_{e}^{{}^{\prime}}$ serves as key and value. The query-to-BEV attention[49],[50] is utilized to propagate object relationships to every pixel in the BEV map. Subsequently, the results are reshaped back into a 2D tensor and passed through a residual block to obtain the refined BEV features $B_{out}^{{}^{\prime}}$. Finally, the easy and hard queries along with $B_{out}^{{}^{\prime}}$ are fed into the progressive decoder .
@@
Figure 7: In the mixing block, queries are used to dynamically aggregate nearby voxel features from each modality. Subsequently, self-attention is employed to aggregate long-range voxel features among instances.
III-E Progressive Decoder
@@
As shown in Fig 2, we decode the BEV features with several deformable attention blocks to obtain the refined queries $q^{{}^{\prime}}$. They are then fed into the (MMVFM) module to decode voxel features $V_{I}^{{}^{\prime}}$ and $V_{L}^{{}^{\prime}}$. As depicted in Fig 7, two separate mixing modules are designed to dynamically aggregate voxel features. In each, we first use several linear layers to regress the proposal box $b_{i}$ from the query $q_{i}^{{}^{\prime}}$. Then, we divide the box into $G=g\times g\times g$ evenly spaced grid points $g_{1},…g_{G}$. Next, we extract grid features $Fg=[f_{g1},…,f_{gG}]$ from $V_{L}$ or $V_{I}$ using voxel pooling. We then embed offsets $\bigtriangleup g_{xk},\bigtriangleup g_{yk},\bigtriangleup g_{zk}$ of each grid point relative to the center of $b_{i}$ into $F_{g}$ through a linear layer, generating shape-aware grid features $F_{g}^{{}^{\prime}}\in\mathbb{R}^{G\times C}$. Following the Adamixer [51] operation, we sequentially perform channel mixing and spatial mixing on them, respectively:
@@
$F_{g}^{{}^{\prime}}={S_{i}}\times transpose(F_{g}^{{}^{\prime}})$ (17)
Here, $C_{i}\in\mathbb{R}^{C\times C}$ and $S_{i}\in\mathbb{R}^{G\times G/4}$ represent the channel and spatial kernels, respectively. After the mixing operation, $F_{g}^{\prime}$ is flattened into a vector and passed through a linear layer to reduce the dimensionality to the final output $F_{I}$ or $F_{L}$. It is worth noting that $S_{i}$ differs in input and output channels from the original design of the adamixer $\mathbb{R}^{G\times G}$, which can reduce the number of parameters by compressing the spatial dimension. Following the aforementioned local mixing operation, we further aggregate long-range 3D features between objects with self-attention mechanisms. Finally, the features from both modalities, along with queries, are concatenated as the output. Subsequently, we decode the fused features through self-attention and a feed-forward layer. Finally, confidences and bounding boxes are regressed by a detection head.
@@ IV EXPERIMENT
Our method is evaluated on the challenging NuScenes dataset. In this section, we first describe the dataset and implementation details, then report the evaluation performance of DDHFusion. Furthermore, we conduct ablation studies and comparison experiments with other homogeneous fusion methods for a thorough analysis.
@@ IV-A Experimental Setup
Implementation Details. Our model is implemented on the MMDetection3D framework. For the LiDAR backbone, we employ VoxelNet [52], while Swin-T serves as the image backbone. The detection range is set to [-54m, 54m] along x and y axes, and [-5m,3m] in the z direction. For ablation studies, the image size is set to $256\times 704$ for faster training. When evaluating the model on the validation or test sets, we follow previous works to set the image size to $384\times 1056$ and double the kernel channels in the LiDAR backbone to achieve better performance. For SAFS, we set score thresholds $d=0.01$, $s=0.25$ according to [41]. The max voxel number $N$ is set to $18000$. DDHFusion is trained on four 4090D GPUs with batch size of $12$. The training process is divided into two stages: (1) we first train the LiDAR-only network for 20 epochs. (2) Then, we combine the image and LiDAR branches and train the fused network for 6 epochs.
During training, we follow CBGS [51] to perform class-balanced sampling and employ the AdamW optimizer with max learning rate $2e^{-3}$. The copy-paste data augmentation [52] is applied to prevent overfitting. We also incorporate the multi-modal copy-paste augmentation[53] in the second stage to stabilize training process.
@@
[…]
PointPillars [54] L 30.5 45.3 68.4 23.0 4.1 23.4 28.8 38.9 27.4 1.1 59.7 30.8
@@
BiCo-Fusion [27] LC 72.4 74.5 88.1 61.9 38.2 73.3 65.7 80.4 78.9 59.8 89.7 88.3
DDHFusion LC

### v2

@@ III-C Homogeneous Voxel Fusion Network
Figure 4: Details of Mamba-based Voxel Fusion. In the left image, for the purpose of visualization, we use a 2D illustration to represent 3D scanning patterns in the HVF. ”Z*2” and ”Z/2” denote doubling or halving the z indices of LiDAR voxels (blue) to align them with image voxels (yellow) or to recover them to their original coordinates. As illustrated in the right figure, we utilize the bidirectional Mamba block in both intra-modal and cross-modal voxel Mamba modules.
Figure 5: Details of Intra-Modal and Cross-Modal BEV Mamba. We apply the four-directional scanning to unfold the BEV feature into 1D sequences, which helps construct a comprehensive spatial relationship. In the cross-modal Mamba block, the parameters for the SS2D modules of each modality are jointly computed from the image and LiDAR BEV features, while $Z_{I}$ and $Z_{L}$ are used to adaptively modulate the weights of their respective modalities.
@@
Where $\bigtriangleup$ denotes the time-scale parameter and $A$ denotes the learnable matrix. $B$, $C$, and $\bigtriangleup$ are inferred from the input sequence $x_{i}$ through linear layers, ensuring the input-specific of the SSM. To ensure complete perception, both forward and backward SSMs are implemented. Their output sequences are added together and modulated by $Y$ to produce the final results.
@@
In CV-Mamba, we directly combine multi-modal voxels in the unified 3D coordinates. It’s important to note that the height range of image voxels $Z_{I}$ is twice that of LiDAR voxels $Z_{L}$. Therefore, we need to double the z indices of LiDAR voxels before combination. Similar to IV-mamba, these voxels are unfolded into a 1D sequence and fed into the SSM. While some voxels may occupy the same position, we treat them as distinct elements in the sequence rather than merging them as in [25]. Then, the output voxels with different modalities are separated and sent back to their respective branches, where they are downsampled by sparse 3D convolution. Notably, in most existing methods, fusing sparse features with different distributions requires operations such as neighbor search [23, 25, 22] or equal-sized grouping [42], while our designed elegant global merge-and-split operation makes feature fusion more straightforward and efficient. Alternately processing voxels with IV-Mamba and CV-Mamba effectively bridges the modality gap and alleviates the negative impact of deteriorated data from each modality.
@@
Finally, inverse sparse convolutions are used to upsample multi-scale voxel features back to the original scale. The output features $V_{I}^{{}^{\prime}}$ and $V_{L}^{{}^{\prime}}$ are not directly combined to avoid interference caused by distinctive distributions. They are fed into the HBF network and the MMVFM module in the decoder.
III-D Homogeneous BEV Fusion Network
@@
In the HBF Network, sparse height compression [43] is utilized to transform $V_{I}^{’}$ and $V_{L}^{’}$ into BEV features. Specifically, It applies max-pooling to voxel features belonging to the same BEV pixel, yielding $B_{I}^{{}^{\prime}}$ and $B_{L}^{{}^{\prime}}$. They are then concatenated with $B_{I}$ and $B_{L}$ to inject more 3D texture and structure details into the BEV space. It is important to note that during SAFS, some potential foreground voxels may be filtered out due to low semantic or depth scores. The LSS compensates for this by retaining image features of the entire environment. Moreover, LSS and SAFS share the same depth map in our work. Since LSS doesn’t undergo feature filtering, it preserves all backpropagated gradients during depth estimation, which indirectly accelerates the convergence of SAFS during training.
@@
Like HVF, we also apply Mamba-based feature fusion including both IB-Mamba and CB-Mamba blocks to generate high-quality BEV features. Firstly, the concatenated features are input into IB-Mamba to globally perceive modality-specific information. Following the SS2D block proposed in [34], we apply the four-directional cross-scanning to flatten BEV tensors, which helps construct a comprehensive spatial relationship. As described in Equations 2 and 3, the SSM is performed simultaneously across all four directions. After passing through LayerNorm, the outputs are summed and modulated by $Y$ for the final result.
@@
In CB-Mamba, to further align multi-modal features in the dense BEV space and adaptively fuse them, we extend the IB-mamba to accommodate dual inputs for cross-modal fusion. As shown in Fig. 5, the parameters for both LiDAR and image SS2D are generated from the concatenated multi-modal BEV features, guiding the feature propagation for both image and LiDAR features. According to the theoretical proof in [34], the SS2D is a special variant of the global transformer. Therefore, CB-Mamba can match corresponding elements between modalities for feature alignment. We describe the SSM parameterization with the following formulas:
@@
Where, the superscript $1-4$ denotes the parameters employed for the four-directional scanning in SS2D. Furthermore, the modulation parameters $Y_{I}$ and $Y_{L}$ are also generated to reweight multi-modal features:
@@
It enables the network to autonomously favor the more reliable modality, thereby improving adaptability in complex scenarios.
@@
After CB-Mamba, the output BEV tensor is fed into the BEV backbone, yielding $B_{out}$, which is then utilized as the input of the PQG module.
@@
Figure 6: Details of Hard Instance Activation. First, self-attention among easy queries is used to extract spatial relationships between objects. Then, query-to-BEV attention is applied to activate latent hard instances within the BEV feature.
III-E Progressive Query Generation
@@
In DDHFusion, the BEV domain is used for efficient generation of target candidates. While many previous works have attempted to achieve this in the range view[44, 45] or sparse 3D space[46, 47], the former is impacted by object occlusions and the latter often relies on center-voting strategies, which are sensitive to uneven voxel or point densities. In contrast, selecting object queries in the BEV domain is relatively simple. First, we generate a heatmap from $B_{out}^{’}$ using a convolutional operation. As described in [14], after processing the heatmap using non-maximum suppression based on 3x3 max pooling, we can obtain queries via the top-k selection. However, there remain some hard targets that tend to be ignored during selection. They often exhibit corrupt information such as sparse surface points or blurred image textures. After height compression, their features may degrade further. Moreover, since targets are typically small in the BEV plane, their features may be easily overshadowed during convolution operations in the BEV backbone.
@@
To address this, [48] propose a multi-stage query generation strategy. It groups queries into different difficulty levels and employs residual blocks to activate BEV features between levels to focus on harder instances. In this work, we simplify the approach by dividing queries into two groups: easy queries $Q_{e}$ and hard queries $Q_{h}$. The corresponding two-stage extraction process is detailed in the following equations:
${H}=Heatmaphead({B_{out}})$ (8)
${P_{e}}=Topk(NMS(H))$ (9)
${Q_{e}}=Collect(B_{out},P_{e})$ (10)
$M=I-Maxpooling(Mask(P_{e}))$ (11)
$B_{out}^{{}^{\prime}}=Resnet({B_{out}})$ (12)
$H^{{}^{\prime}}=Heatmaphead({B_{out}^{{}^{\prime}}})$ (13)
$P_{h}=Topk(NMS(M\cdot{H^{{}^{\prime}}}))$ (14)
${Q_{h}}=Collect(B_{out}^{{}^{\prime}},P_{h})$ (15)
@@
Where $P_{e}$ and $P_{h}$ denote the positions of easy and hard queries. $M$ represents the mask generated from $P_{e}$, which is used to obtain $P_{h}$ by masking the expanded regions corresponding to $Q_{e}$. During training, we need to compute the heatmap loss of the second stage that doesn’t lie within the masked regions of $Q_{e}$, guiding the network to focus on hard targets. However, activating BEV feature by the residual block has limited effectiveness—it can only aggregate local information from difficult regions to enhance responses. In the driving environment, the spatial relationships between objects exhibit certain regularities, such as the placement of vehicles and barriers. Based on this observation, we propose a hard instance activation (HIA) module, which leverages easy queries $Q_{e}$ to activate features in difficult regions. Specifically, we modify Equation 12 as follows:
${B_{out}^{{}^{\prime}}}=HIA({\color[rgb]{1,0,0}Q_{e}},{B_{out}})$ (16)
@@
As shown in Fig 6, in HIA, we first embed categories – in form of one-hot vectors – and positions into the easy queries. Then, we utilize self-attention to extract relationships between objects, resulting in enhanced queries $Q_{e}^{{}^{\prime}}$. Then, $B_{out}$ is flattened and regarded as the query and $Q_{e}^{{}^{\prime}}$ serves as the key and value. The query-to-BEV attention[49],[50] is utilized to propagate object relationships to every pixel in the BEV map. Subsequently, the results are reshaped back into a 2D tensor and passed through a residual block to obtain the refined BEV features $B_{out}^{{}^{\prime}}$ for the selection of hard queries $Q_{h}$. Finally, the easy and hard queries along with $B_{out}^{{}^{\prime}}$ are fed into the progressive decoder .
@@
Figure 7: In the mixing block, queries are used to dynamically aggregate nearby voxel features from each modality. Subsequently, self-attention is employed to extract long-range features among instances.
III-F Progressive Decoder
@@
As shown in Fig 2, we decode the BEV features with several deformable attention blocks to obtain the refined queries $q^{{}^{\prime}}$. They are then fed into the (MMVFM) module to decode voxel features $V_{I}^{{}^{\prime}}$ and $V_{L}^{{}^{\prime}}$. As depicted in Fig 7, two separate mixing modules are designed to dynamically aggregate voxel features. In each, we use several MLPs to regress the proposal box $b_{i}$ from the query $q_{i}^{{}^{\prime}}$. $b_{i}$ is evenly divided into $G=g\times g\times g$ grid points $g_{1},…g_{G}$. Next, we extract grid features $Fg=[f_{1},…,f_{G}]$ from $V_{L}$ or $V_{I}$ using voxel pooling. We then embed offsets $\bigtriangleup g_{xk},\bigtriangleup g_{yk},\bigtriangleup g_{zk}$ of each grid point relative to the center of $b_{i}$ into $F_{g}$ through a linear layer, generating shape-aware grid features $F_{g}^{{}^{\prime}}\in\mathbb{R}^{G\times C}$. Following Adamixer [51], we sequentially perform channel mixing and spatial mixing on them:
@@
$F_{g}^{{}^{\prime\prime}}=transpose(F_{g}^{{}^{\prime}})\times{S_{i}}$ (20)
Here, $C_{i}\in\mathbb{R}^{C\times C}$ and $S_{i}\in\mathbb{R}^{G\times G/4}$ represent the channel and spatial kernels, respectively. After the mixing operation, $F_{g}^{\prime\prime}$ is flattened into a vector and passed through a linear layer to reduce the dimensionality, resulting in the output $F_{I}$ or $F_{L}$. It is worth noting that $S_{i}$ differs in input and output channels from the original design of Adamixer $\mathbb{R}^{G\times G}$. Ours can reduce the number of parameters by compressing the spatial dimension. Following the aforementioned local mixing operation, we further aggregate long-range 3D features between objects with self-attention. Finally, the features from both modalities, along with queries, are concatenated as the output. Subsequently, we decode the output features through self-attention and a feed-forward layer. Finally, a detection head is applied to regress confidences and bounding boxes.
@@ IV EXPERIMENT
Our method is evaluated on the challenging NuScenes dataset. In this section, we first describe the dataset and implementation details, then report the evaluation performance of DDHFusion. Furthermore, we conduct ablation studies of each component and comparison experiments with other homogeneous fusion methods for a thorough analysis.
@@ IV-A Experimental Setup
Method Modality mAP NDS Car Truck C.V. Trailer Bus Barrier Motor. Bike Ped. T.C.
@@
BiCo-Fusion [27] LC 72.4 74.5 88.1 61.9 38.2 65.7 73.3 80.4 78.9 59.8 89.7 88.3
DDHFusion(ours) LC 73.8 75.6 90.4 63.6 39.1 66.7 74.4 81.0 81.8 63.6 90.3 87.5
Implementation Details. Our model is implemented on the MMDetection3D framework. For the LiDAR backbone, we employ VoxelNet [58], while Swin-T serves as the image backbone. The detection range is set to $[-54m,54m]$ along x and y axes, and $[-5m,3m]$ in the z direction. For ablation studies and extensive comparative experiments, the image size is set to $256\times 704$ for fast training. When evaluating the model on the validation or test sets, we follow previous works to set the image size to $384\times 1056$ and double the kernel channels in the LiDAR backbone to achieve better performance. For SAFS, we set score thresholds $d=0.01$, $s=0.25$. The maximum voxel number $N$ is set to $18000$. DDHFusion is trained on four NVIDIA 4090D GPUs with batch size of $12$. The training process is divided into two stages: (1) training the LiDAR-only network for 20 epochs. (2) combining the image and LiDAR branches and training the fused network for 6 epochs.
During training, we follow CBGS [51] to perform class-balanced sampling and employ the AdamW optimizer with maximum learning rate $2e^{-3}$. The copy-paste data augmentation [58] is applied to prevent overfitting.